A circuit consists of a 292- resistor and a 0.206-H inductor. These two elements are connected in series across a generator that has a frequency of 93.6 Hz and a voltage of 251 V. (a) What is the current in the circuit? (b) Determine the phase angle between the current and the voltage of the generator. Note: The ac current and voltage are rms values and power is an average value unless indicated otherwise.
Answer #1

Solution)

Given,

resistor R = 292 Ω

inductor L = 0.206 H

frequency f = 93.6 Hz

voltage V = 251 V

(a) Impedance in the circuit

Z = √ R^2 +XL^2

= √ (292)^2 + (2πf L)^2

= √ [ (85264) + 14662.4

= 316.11 Ω

current I = V/Z = 251/ 316.11

= 0.794 A (Ans)

======

(b) Phase angle

φ = tan^-1( XL/R)

= 22.52 deg (Ans)

======

Good luck!:)